OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 3x3 + 27x2y - 24xy - 216y2 and check your factors for accuracy. If someone could please explain this to me, that would be great because I'd like to be able to learn how to do this on my own. :)

OpenStudy (jdoe0001):

let's first group it in 2 pairs, \(\bf (3x^3 + 27x^2y) - (24xy + 216y^2)\) can you get common factors for each of those pairs?

OpenStudy (anonymous):

3x3+27x2y-24xy-216y2= 3x2(x+9y)-24y(x+9y)= (3x2-24y) (x+9y)= 3(x2-8y) (x+9y)

OpenStudy (anonymous):

For the first pair, would the common factor be three?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

well, a bit more than 3

OpenStudy (anonymous):

Okay, so would I have to divide the 27x2y by three?

OpenStudy (uri):

3x3 + 27x2y - 24xy - 216y2 First pair will be 3x^2(3+9y) and second will be -24(x+9y) So, (x+9y)(3x^2-24)

OpenStudy (uri):

-24y*

OpenStudy (anonymous):

I don't get it, I need an explanation, like how do you find those numbers? :(

OpenStudy (uri):

Hm what's common in 3x^3 +27x^2y? <--- 3 x 1= 3 and 3 x 9=27

OpenStudy (uri):

Get it till here?

OpenStudy (anonymous):

I understand that 3 and 27 have three in common but where does the (3+9y) come from? Because the 27 has an x^2 as well....

OpenStudy (jdoe0001):

$$ 3x^3 + 27x^2y - 24xy - 216y^2 \implies (3x^3 + 27x^2y) - (24xy + 216y^2)\\ \color{green}{\text{now we get common factors on both}}\\ 3x^2\color{blue}{(x + 9y)} - 24y\color{blue}{(x + 9y)}\\ \color{green}{\text{now we see 2 common factors and we take them out}}\\ (3x^2-24y)(x + 9y)\\ $$

OpenStudy (anonymous):

I don't understand.... :(

OpenStudy (uri):

@Breeziee Try asking your teacher?

OpenStudy (anonymous):

She isn't online right now. I just don't understand how you get the x+9y thing is that by dividing the 27 thing by three

OpenStudy (uri):

We take the common from 3x^3+27x^2y 3x^2(x+9y) .....3x^2 is common!!

OpenStudy (jdoe0001):

\(\bf (3x^3 + 27x^2y) \implies (3xxx + 3 \times 9xxy)\) common factor, see what's common to both?

OpenStudy (anonymous):

Yes, I see it now

OpenStudy (uri):

Thankgod :P

OpenStudy (anonymous):

Sorry, I'm just dumb. :( So what about the second half of the equation?

OpenStudy (uri):

24xy - 216y2 24 x1 =24 and 24x9 =126 In this equation 24xy has one y and 216 has 2 y's how many are common? 1 y So 24y(x+9y)

OpenStudy (uri):

216* not 126..sorry typo.

OpenStudy (anonymous):

Okay, I get that. Now what's next?

OpenStudy (jdoe0001):

\(\bf (24xy + 216y^2) \implies (24xy + 24\times 24yy)\) common factor there too

OpenStudy (jdoe0001):

\(3x^3 + 27x^2y - 24xy - 216y^2 \implies (3x^3 + 27x^2y) - (24xy + 216y^2)\\ \color{green}{\text{now we get common factors on both}}\\ 3x^2\color{blue}{(x + 9y)} - 24y\color{blue}{(x + 9y)}\\ \color{green}{\text{now we see 2 common factors and we take them out}}\\ (3x^2-24y)(x + 9y)\)

OpenStudy (uri):

The just that, 3x^2(x+9y) -24y(x+9y) (x+9y) is same so we write it only one time and 3x^2-24y (x+9y)(3x^2-24y)

OpenStudy (anonymous):

Oh, thank you so much, I think I get it now!

OpenStudy (uri):

Yw :)

OpenStudy (anonymous):

Thank you so much though for putting up with me, I can be a little difficult sometimes. :)